# Exploring the parameter dependence of atomic minima with implicit differentiation

### v1

@@ 3.2 Numerical details on LAMMPS constraint minimization
For $\alpha_{0}<10^{-7}$, the error is large due to limitations in numerical precision. For larger $\alpha_{0}$ values, the error does not change. However, the minimization time increases significantly for $\alpha_{0}\geq 10^{-3}$. Therefore, we conclude that values of $\alpha_{0}\in[10^{-6};10^{-4}]$ are optimal for the energy method. For the constraint energy minimization, we use the FIRE algorithm implemented in LAMMPS.
@@ 4.1 Adaptive step for loss minimization
This section describes the adaptive step calculation for the inverse design applications. As explained in the main text (Results H), at iteration $k+1$ of the minimization procedure, the potential parameters are updated as ${{\bm{\Theta}}}^{(k+1)}={{\bm{\Theta}}}^{(k)}-h\nabla_{{{\bm{\Theta}}}}L({{\bm{\Theta}}}^{(k)})$ and positions as ${\bf X}^{*}_{{{\bm{\Theta}}}^{(k+1)}}={\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}+({{\bm{\Theta}}}^{(k+1)}-{{\bm{\Theta}}}^{(k)})\nabla_{{{\bm{\Theta}}}}{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}$. Accordingly, the loss at iteration $k+1$ is
$L({{\bm{\Theta}}}^{(k+1)})=\frac{1}{2}\|{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}+h\Delta{\bf X}^{(0)}_{{{\bm{\Theta}}}^{(k)}}-{\bf X}^{*}\|^{2},$ (40)
where $\Delta{\bf X}^{(0)}_{{{\bm{\Theta}}}^{(k)}}$ is the change in atomic positions with step $h=1$. Then, the change in loss at a given iteration is
$\Delta L({{\bm{\Theta}}}^{(k+1)})\equiv L({{\bm{\Theta}}}^{(k+1)})-L({{\bm{\Theta}}}^{(k)})=h\Delta{\bf X}^{(0)\,\top}_{{{\bm{\Theta}}}^{(k)}}\big{(}{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}-{\bf X}^{*}\big{)}+\frac{1}{2}h^{2}\|\Delta{\bf X}^{(0)}_{{{\bm{\Theta}}}^{(k)}}\|^{2}.$ (41)
@@
$h(k)=-\frac{\Delta{\bf X}^{(0)\,\top}_{{{\bm{\Theta}}}^{(k)}}\big{(}{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}-{\bf X}^{*}\big{)}}{\|\Delta{\bf X}^{(0)}_{{{\bm{\Theta}}}^{(k)}}\|^{2}}.$ (42)
@@ 4.2 W-Be POTENTIAL FINE-TUNING
Figure 9 presents the error minimization for the potential fine-tuning for the W-Be system presented in the main text (Results H).
Figure 9: Loss minimization for W-Be potential fine-tuning problem. a Implicit loss during the minimization procedure as defined in equation (9) of the main text. b Difference between a current and target Be atom positions during minimization.

### v2

@@ 3.2 Numerical details of LAMMPS constraint energy minimization
For $\alpha_{0}<10^{-7}$, the error is large due to limitations in machine precision. For larger $\alpha_{0}$ values, the error does not change. However, the number of energy minimization iterations increases significantly for $\alpha_{0}\geq 10^{-3}$. Therefore, we conclude that values of $\alpha_{0}\in[10^{-6};10^{-4}]$ are optimal for the energy implicit derivative method. For the constraint energy minimization, we use the FIRE algorithm implemented in LAMMPS.
@@ 4.1 Adaptive step for loss minimization
This section describes the adaptive step calculation for the inverse design applications. As explained in the main text (section IIIH), at iteration $k+1$ of the minimization procedure, the potential parameters are updated as ${{\bm{\Theta}}}^{(k+1)}={{\bm{\Theta}}}^{(k)}-h\nabla_{{{\bm{\Theta}}}}L({{\bm{\Theta}}}^{(k)})$ and positions as ${\bf X}^{*}_{{{\bm{\Theta}}}^{(k+1)}}={\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}+({{\bm{\Theta}}}^{(k+1)}-{{\bm{\Theta}}}^{(k)})\nabla_{{{\bm{\Theta}}}}{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}={\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}-h\nabla_{{{\bm{\Theta}}}}L({{\bm{\Theta}}}^{(k)})\nabla_{{{\bm{\Theta}}}}{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}$. Accordingly, the loss at iteration $k+1$ is
$L({{\bm{\Theta}}}^{(k+1)})=\frac{1}{2}\|{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}+h\Delta{\bf X}^{(0)}_{{{\bm{\Theta}}}^{(k)}}-{\bf X}^{*}_{\rm ref}\|^{2},$ (41)
where $\Delta{\bf X}^{(0)}_{{{\bm{\Theta}}}^{(k)}}$ is defined as the change in atomic positions corresponding to step $h=1$, $\Delta{\bf X}^{(0)}_{{{\bm{\Theta}}}^{(k)}}\equiv-\nabla_{{{\bm{\Theta}}}}L({{\bm{\Theta}}}^{(k)})\nabla_{{{\bm{\Theta}}}}{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}$. Then, the change in loss at a given iteration is
$\Delta L({{\bm{\Theta}}}^{(k+1)})\equiv L({{\bm{\Theta}}}^{(k+1)})-L({{\bm{\Theta}}}^{(k)})=h\Delta{\bf X}^{(0)\,\top}_{{{\bm{\Theta}}}^{(k)}}\big{(}{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}-{\bf X}^{*}_{\rm ref}\big{)}+\frac{1}{2}h^{2}\|\Delta{\bf X}^{(0)}_{{{\bm{\Theta}}}^{(k)}}\|^{2}.$ (42)
@@
$h(k)=-\frac{\Delta{\bf X}^{(0)\,\top}_{{{\bm{\Theta}}}^{(k)}}\big{(}{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}-{\bf X}^{*}_{\rm ref}\big{)}}{\|\Delta{\bf X}^{(0)}_{{{\bm{\Theta}}}^{(k)}}\|^{2}}.$ (43)
Accordingly, the atomic position change at iteration $k+1$ is
$\Delta{\bf X}_{{{\bm{\Theta}}}^{(k)}}=h(k)\nabla_{{\bm{\Theta}}}L({{\bm{\Theta}}}^{(k)})\nabla_{{\bm{\Theta}}}{\bf X}^{*}_{{{\bm{\Theta}}}^{(k)}}.$ (44)
@@ 4.2 W-Be Potential fine-tuning
Figure 9 presents the loss minimization for the potential fine-tuning for the W-Be system presented in the main text (section IIH).
Figure 9: Loss minimization for W-Be potential fine-tuning. a Implicit loss during the minimization procedure as defined in equation (9) of the main text. b Difference between a current and target Be atom positions during minimization.